3) A $1,060 face value bond is selling in the market place for $927. It matures in 3 years. If keep to maturity, what is the bond's yield to maturity? (Round to the nearest thousands and convert to a percentage).
Answer #1

Market price = Face value / (1+r)n

Market price = 927

Face value = 1060

r = yield to maturity

n = 3

927 = 1060/(1+r)3

(1+r)3=(1060/927)

(1+r) = (1060/927)(1/3​​​​​)

1+r = 1.046

r = 0.046

r = 4.6%